OpenStudy (anonymous):

Two pipe running together can fill a cistern in 3.5mins and one pipe takes 3 mins more than the other find the time in which each pipe would fill cistern

OpenStudy (anonymous):

@Michele_Laino @pooja195 @perl

OpenStudy (perl):

Let x be the rate of the first pipe and y is the rate of the second pipe. And t be the time it takes for first pipe to fill cistern. Then we have the following equations, use d = r* t 1 cistern = (x+y) * 3.5 min 1 cistern = x * t 1 cistern = y * (t + 3)

OpenStudy (anonymous):

I can't understand the above.

OpenStudy (perl):

Have you used the formula distance = rate * time

OpenStudy (anonymous):

Yes..

OpenStudy (perl):

I am using a similar equation. 1 cistern = rate of pipe * time

OpenStudy (anonymous):

What next?

OpenStudy (perl):

Let x be the rate of the first pipe and y is the rate of the second pipe. And t be the time it takes for first pipe to fill cistern. 1 cistern = (x+y) * 3.5 min 1 cistern = x * t 1 cistern = y * (t + 3)

OpenStudy (anonymous):

3 cistern?

OpenStudy (perl):

they are three different equations

OpenStudy (perl):

"Two pipe running together can fill a cistern in 3.5mins" That gives us 1 cistern = (x+y) * 3.5 min , where x and y are the rates of the pipes

OpenStudy (anonymous):

Help PLEASE?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I think that @perl gave you the right explanation

OpenStudy (michele_laino):

nevertheless I can give you another way to solve your problem. The working rates of the two pipes are: \[\Large\frac{W}{x},\quad \frac{W}{{x + 3}}\]

OpenStudy (michele_laino):

where W is the work to be done

OpenStudy (michele_laino):

now, following the text of your problem, we can write: \[\Large \frac{W}{x} + \frac{W}{{x + 3}} = \frac{W}{{3.5}}\] and, simplifying that expression for W, we get: \[\Large \frac{1}{x} + \frac{1}{{x + 3}} = \frac{1}{{3.5}}\] Please solve that equation for x

OpenStudy (anonymous):

x+3+x/x^2+3x = 1/3.5?

OpenStudy (michele_laino):

first step: we have: \[\Large \frac{1}{x} + \frac{1}{{x + 3}} = \frac{2}{7}\]

OpenStudy (michele_laino):

yes! since 3.5= 7/2

OpenStudy (michele_laino):

now, the least common multiple, of x, x+3 and 7, is: x*(x+3)*7 am I right?

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

is that equation is equivalent to this one? \[\Large 1 \times 7 \times \left( {x + 3} \right) + 7x = 2x\left( {x + 3} \right)\]

OpenStudy (michele_laino):

oops..is that equation equivalent to this one?

OpenStudy (anonymous):

I guess..

OpenStudy (michele_laino):

more steps: \[\Large \frac{{1 \times 7\left( {x + 3} \right)}}{{7x\left( {x + 3} \right)}} + \frac{{1 \times 7x}}{{7x\left( {x + 3} \right)}} = \frac{{2x\left( {x + 3} \right)}}{{7x\left( {x + 3} \right)}}\]

OpenStudy (michele_laino):

now?

OpenStudy (anonymous):

I'm lost...

OpenStudy (michele_laino):

as denominator, of each fraction, we find the least common multiple

OpenStudy (anonymous):

Getting it.. Now?

OpenStudy (michele_laino):

now, if we consider the first fraction, for example, we have: